For the given ee value, calculate the percentage o
Answer #1

obtained ee of compound = 91 %

percentage of Enantiomer A = 91 + 9 / 2 = 95.5

Percentage of Enantiomer B = 9/2 = 4.5

Be sure to answer all parts. For the given ee value, calculate the percentage of each enantiomer present. Assume A is in exce